OpenStudy (theeric):

Hello! I'm working on an integral for part of a problem, where I have to solve for \(B\). It looks something like this: \(\Large\varepsilon=\int_0^\infty(Be^{cx}\mp1)^{-1}dx\) This would be so much easier if the integrand was to the first power! Thanks for any help.

OpenStudy (theeric):

Later I will want \(\large e^{-cx}\) still, so that the integral converges.

OpenStudy (anonymous):

If it weren't for that -/+ the integral could be done by u-sub, which would lead to partial fractions and it could be done from there. Not sure how any of those methods change when you throw that -/+ in there O.o

OpenStudy (theeric):

Well, I'm just using that to combine two equations for different physical situations. They are \(\Large\varepsilon=\int_0^\infty(Be^{cx}-1)^{-1}dx\) and \(\Large\varepsilon=\int_0^\infty(Be^{cx}+1)^{-1}dx\) I've been looking into substitutions, but haven't come up with anything helpful yet, I think. I've considered \(\large u=Be^{cx}+1\), for example. But this leads to... Oh... Hmm... Hehe, I should take another look at that! Thanks!

OpenStudy (anonymous):

Yeah do that u-sub, it'll lead to a partial fractions problem and it comes up very easily.

OpenStudy (theeric):

I'll give that a shot...

OpenStudy (anonymous):

Turning out okay? :)

OpenStudy (theeric):

\(\large{ u=Be^{cx}\mp1\\\quad\Rightarrow x=\dfrac{\ln\left(\dfrac{u\pm1}B\right)}c\\\quad\Rightarrow \text du=Bce^{cx}\text dx\\\quad\quad\Rightarrow dx=\left(Bce^{cx}\right)^{-1}du}\) This looks ugly so far..

OpenStudy (anonymous):

Nuu, no need to do that substitution

OpenStudy (anonymous):

Ill start ya out :P

OpenStudy (theeric):

For \(x\)? \(\text dx\) depends on \(x\) though, right? I'll watch what you do.

OpenStudy (anonymous):

\[u = Be^{cx} + 1\] \[\frac{ u-1 }{ B } = e^{cx}\] \[du = Bce^{cx}dx\] \[dx = \frac{ du }{ Bce^{cx} }\] All of this gives us: \[\int\limits_{0}^{\infty}\frac{ 1 }{ Bce^{cx}*u }du = \int\limits_{0}^{\infty}\frac{ 1 }{ Bc \frac{ u-1 }{ B } *u}du = \frac{ 1 }{c } \int\limits_{0}^{\infty}\frac{ 1 }{ u(u-1) }du\] So the key was to solve for e^(cx), not x itself.

OpenStudy (theeric):

Oh, thank you very much! And now you suggest partial fractions, right? Otherwise I would tackle this with integration by parts just because I'm more familiar with that. But I have been taught to use partial fractions.

OpenStudy (anonymous):

Yeah, partial fractions seems more natural to me. I see factoring on bottom, so that keys to me partials. If you know how to go about that, it should come out pretty well

OpenStudy (theeric):

Okay, thanks! I'll work on that, then.

OpenStudy (anonymous):

Alright, cool. I gotta head out now, so good luck :) Glad that helps.

OpenStudy (theeric):

Thank you for your help! Take care! :)

OpenStudy (theeric):

We left off with \[\large{\dfrac1c\int_0^\infty \dfrac1{u(u\pm1)}du}\] This becomes \[\large{\dfrac1c}\int_0^\infty \left(\dfrac {A(u\pm1)+Bu}{u(u\pm1)}\right)du\]\(\qquad\)Note, \(1=A(u\pm1)+Bu\) \(\qquad\qquad\)So, \(u=0\Rightarrow1=\pm A\qquad\Rightarrow\qquad A=\mp1\) \(\qquad\qquad\) and \(u=\mp1\Rightarrow1=\mp B\qquad\Rightarrow\qquad B=\pm1\) \[\large{\dfrac1c}\int_0^\infty \left(\dfrac Au+\dfrac B{u\pm1}\right)du\] Putting this together, we have \[\large{\dfrac1c}\int_0^\infty \left(\dfrac {\mp1}u+\dfrac {\pm1}{u\pm1}\right)du\] So, there is the partial fractions part.

OpenStudy (theeric):

I made a sign error when determining \(A\) and \(B\). \(\large u=0\Rightarrow1=\pm A\qquad\Rightarrow\qquad A=\pm1\) \(u=\mp1\Rightarrow1=\mp B\qquad\Rightarrow\qquad B=\mp1\) There was no justification for this mistake. We can break the integral up to be \[\large{\dfrac1c}\left(\int_0^\infty \dfrac {\pm1}u\text du+\int_0^\infty\dfrac {\mp1}{u\pm1}\text du\right)\] It can be made to look nicer, like \[\large{\dfrac1c}\left(\pm\int_0^\infty u^{-1}\text du\mp\int_0^\infty(u\pm1)^{-1}\text du\right)\] This becomes \[\large{\dfrac1c}\left(\pm\left[\ln(u)\right]_0^\infty\mp\left[\ln(u\pm1)\right]_0^\infty\right)\] Now we go back to \(\large u=Be^{cx}\mp1\). With substitution, we have \[\large{\dfrac1c}\left(\pm\left[\ln(Be^{cx}\mp1)\right]_0^\infty\mp\left[\ln(Be^{cx}\mp1\pm1)\right]_0^\infty\right)\]which simplifies to be \[\large{\dfrac1c}\left(\pm\left[\ln(Be^{cx}\mp1)\right]_0^\infty\mp\left[\ln(Be^{cx})\right]_0^\infty\right)\] For clarity, the last expression captures two: \[\large{\dfrac1c}\left(+\left[\ln(Be^{cx}-1)\right]_0^\infty-\left[\ln(Be^{cx})\right]_0^\infty\right)\]and \[\large{\dfrac1c}\left(-\left[\ln(Be^{cx}+1)\right]_0^\infty+\left[\ln(Be^{cx})\right]_0^\infty\right)\]

OpenStudy (theeric):

Another mistake I noticed is that I left the bounds the same when integrating with respect to \(u\). I should've noted that they were bounds with respect to \(x\), or omitted them altogether. I think that I can simpify like this:\[\large{\dfrac1c}\left(\pm\left[\ln(Be^{cx}\mp1)\mp\ln(Be^{cx})\right]_0^\infty\right)\] This allows the natural log functions to be multiplied or divided (depending on sign), \[\large{\dfrac1c}\left(\pm\left[\ln\left((Be^{cx}\mp1)(Be^{cx})^{\mp1}\right)\right]_0^\infty\right)\] Now it will be better to work with two different equations. Distributing with the top operator we get \[\large{\dfrac1c}\left(+\left[\ln\left((Be^{cx}-1)(Be^{cx})^{-1}\right)\right]_0^\infty\right)\\~\\=\large{\dfrac1c}\left(\left[\ln\left(1-(Be^{cx})^{-1}\right)\right]_0^\infty\right) \\~\\=\large{\dfrac1c}\left(\left[\ln\left(1-B^{-1}e^{-cx}\right)\right]_0^\infty\right) \\~\\=\large{\dfrac1c\left(\ln(1)-\ln(1-B^{-1})\right)} \\~\\=\large{\dfrac1c\ln\left(\dfrac1{1-\dfrac1B}\right)=\varepsilon} \] Okay! \[\large{\ln\left(\dfrac1{1-\dfrac1B}\right)=c\varepsilon \\~\\\Rightarrow \dfrac1{1-\dfrac1B}=e^{c\varepsilon} \\~\\\Rightarrow 1-\dfrac1B=e^{-c\varepsilon} \\~\\\Rightarrow\dfrac1B=1-e^{-c\varepsilon} \\~\\\Rightarrow B=(1-e^{-c\varepsilon})^{-1}}\] And, if I go through with the bottom operations, it should come out to \(B=-(1-e^{-c\varepsilon})^{-1}\).

OpenStudy (theeric):

For the sake of anyone reading this, it seems that I made a mistake. There should be no exponent of \(\mp1\), it should just be \(-1\). Changing that, I got to the correct result. Also, my book provides a formula for this that differs in that exponent only. I haven't found my mistake, but I might search for it another time.